A firm has 3,490,000 authorized shares, 520,000 issued shares, and 340,000 outstanding shares before completing a stock split. The shares have a par value of $0.09 before the stock split. The firm completes a reverse 2-to-1 split. What is the value in the common stock account after the stock split?
Answer #1

Total shares issued = 520,000

Total shares outstanding = 340,000

Shares Reverse split as 2-to-1.

Total shares after split = 340,000/2 = 170,000

Value of shares after reverse split = 170,000 * 0.09 = 15,300

Remaining shares = 520,000-340,000 = 180,000

Value of these shares = 180,000 * 0.09 = 16,200

Total value in the common stock accout after stock split = 16,200+15,300 = 31,500.